OpenStudy (anonymous):

vanillin is composed of carbon,63.2%; hydrogen,5.26% and oxygen, 31.6%. What is the empirical formula of vanillin? I tried changing the percentages into grams then converted them to moles, after that I divided each element by the smallest amount of moles then multiplied to get whole numbers but my answer was wrong. Step by step work out please. thx :)

OpenStudy (anonymous):

ok so you did well first step: C - 63,2/12,011 = 5,26184 H - 5,26/1,008 = 5,218254 O - 31,6/15,999=1,975123 now you divide with O C - 5,26184/1,975123 = 2,664 H - 5,218254/1,975123= 2,642 O - 1,975123/1,975123 = 1 now you have their atomic ratios and now you need to guess, i mean multiply, lets try with 2 2*2,664 = 5,328 and right away you know it isnt 2, so lets try 3 3*2,664 = 7,992 3*2,642 = 7,926 3*1 = 3 so that is your answer C8H8O3

OpenStudy (anonymous):

thank you!:)